# Degeneration of families of projective hypersurfaces and Hodge conjecture

### v10

@@ 3 Hodge conjecture for smooth projective varieties
$f:\mathcal{Y}^{\prime}=V(\tilde{f})\subset\mathbb{P}^{d+1}_{\mathbb{C}}\times\mathbb{P}(l_{X})\to\mathbb{P}(l_{X})$
the universal family of projective hypersurfaces of degree $l_{X}$, and $0_{X}\in\mathbb{P}(l_{X})$ the point corresponding to $Y_{0_{X}}$. Denote $\Delta\subset\mathbb{P}(l_{X})$ the discriminant locus parametrizing the hypersurfaces which are singular. Let $\bar{S}:=\bar{0_{X}}^{\mathbb{Q}}\subset\mathbb{P}(l_{X})$ be the $\mathbb{Q}$-Zariski closure of $0_{X}$ inside $\mathbb{P}(l_{X})$. If $Y_{0_{X}}$ is smooth, then $X$ satisfy the hodge conjecture since $Y_{0_{X}}$ satisfy the Hodge conjecture by [5] and by the induction hypothesis since $X$ is birational to $Y_{0_{X}}$ by proposition 1(ii). Hence, we may assume that $Y_{0_{X}}$ is singular, that is $0_{X}\in\Delta$. Then, $\bar{S}\subset\Delta$ since $\Delta\subset\mathbb{P}(l_{X})$ is defined over $\mathbb{Q}$. Let $\bar{T}\subset\mathbb{P}(l_{X})$ be an irreducible closed subvariety of dimension $\mathop{\rm dim}\nolimits(\bar{S})+1$ defined over $\mathbb{Q}$ such that
@@
We take for $\bar{T}=V(f_{1},\cdots,f_{s})\subset\mathbb{P}(l_{X})$ an irreducible complete intersection which intersect $\Delta$ properly and such that
  • $V(f_{1})\supset\mathop{\rm sing}\nolimits(\Delta)\cup S$
  • for each $r\in[1,\cdots,s]$, $V(f_{1},\cdots,f_{r})\supset\mathop{\rm sing}\nolimits(V(f_{1},\cdots,f_{r-1})\cap\Delta)\cup S$ contain inductively the singular locus of the intersection with $\Delta$ and $S$.
@@
is an isomorphism if $\mathop{\rm dim}\nolimits\bar{T}\geq 2$ and surjective if $\mathop{\rm dim}\nolimits\bar{T}=1$. We have then $\bar{T}\cap\Delta=\bar{S}\cup\bar{S}^{\prime}$ where $\bar{S}^{\prime}\subset\Delta$ is a divisor. We then consider the family of algebraic varieties defined over $\mathbb{Q}$
@@
  • $t\in S(\mathbb{C})$ is the point such that $\mathcal{Y}^{\prime}_{\mathbb{C},t}=Y_{0_{X}}$ and $\epsilon_{t}:E_{0}:=\mathcal{E}_{0,\mathbb{C},t}\to Y_{0_{X}}$ is surjective, in particular $E_{0}$ is birational to $Y_{0_{X}}$, hence $E_{0}$ is birational to $X$. We denote $E_{i}:=\mathcal{E}_{i,\mathbb{C},t}$ for each $0\leq i\leq r$ and $E:=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$.
@@
which is an algebraic variety by [6]. Note that $\mathbb{V}^{p}_{S}(\lambda)\subset HL^{p,2p}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$ is the union of the irreducible components passing through $\lambda$. Similarly, for $\eta\in H^{2p}(\mathcal{E}_{\mathbb{C},t^{\prime}}^{an},\mathbb{Q})$, $t^{\prime}\in S(\mathbb{C})$, we consider $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$ the universal covering in $\mathop{\rm AnSm}\nolimits(\mathbb{C})$ and
@@
$(E^{2p}_{DR}(\mathcal{Y}_{T^{o}}/T^{o}),F)=H^{2p}\tilde{f}_{X*}E_{zar}(\Omega^{\bullet}_{\mathcal{Y}_{T^{o}}/T^{o}})\in\mathop{\rm Vect}\nolimits_{\mathcal{D}fil}(T^{o})$
@@
$(E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o}}/T^{o}),F):=H^{2p}\tilde{f}_{X*}E_{zar}(\Omega^{\bullet}_{\mathcal{Y}_{T^{o}}/T^{o}}(\mathop{\rm log}\nolimits\mathcal{E}))\in\mathop{\rm Vect}\nolimits_{\mathcal{D}fil}(\bar{T}^{o})$
@@
$\alpha\in H^{2p}(\mathcal{Y}_{t^{\prime\prime}}^{an},\mathbb{Q})=\mathbb{H}^{2p}(\mathcal{E}^{an}_{t^{\prime}},\psi_{S}R\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}_{T^{o}}}),\,t^{\prime}\in S_{\mathbb{C}},t^{\prime\prime}\in T^{o}(\mathbb{C}),$
@@
  • $\pi_{T^{o}}:\tilde{T}^{o,an}\to T_{\mathbb{C}}^{o,an}$ the universal covering in $\mathop{\rm AnSm}\nolimits(\mathbb{C})$, the factorization $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to\tilde{T}^{o,an}\xrightarrow{\pi_{T^{o}}}T_{\mathbb{C}}^{o,an}$ and
@@
    which is an algebraic subvariety by [6],
@@
    $\mathbb{V}_{S}(\alpha):=\pi_{S}(\alpha\times\tilde{S}^{an}_{\mathbb{C}})\subset H^{2p}\psi_{S}E_{DR}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})^{an}$
@@
where $HL^{2p,p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$, $HL^{2p,p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})$ are the locus of Hodge classes. Let $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$, where we recall $E_{0}=\mathcal{E}_{0,\mathbb{C},t}$. Consider then $\tilde{\lambda}\in H^{2p}(E^{an},\mathbb{Q})$, such that $q(\tilde{\lambda})=\lambda$, and
$\tilde{\lambda}:=Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\in Sp_{E/f_{X}}(H^{2p}(E^{an},\mathbb{Q})).$
@@
using the fact that there exists a neighborhood $V_{S\cup S^{\prime}}\subset\bar{T}_{\mathbb{C}}$ of $S_{\mathbb{C}}\cup S^{\prime}_{\mathbb{C}}$ in $\bar{T}_{\mathbb{C}}$ for the usual complex topology such that the inclusion $i_{\mathcal{E}^{\prime}}:\mathcal{E}^{\prime}_{\mathbb{C}}\hookrightarrow\mathcal{Y}_{\mathbb{C},V_{S\cup S^{\prime}}}$ retraction $r:\mathcal{Y}_{\mathbb{C},V_{S\cup S^{\prime}}}\to\mathcal{E}^{\prime}_{\mathbb{C}}$ which is an homotopy equivalence. The locus of Hodge classes passing through $\lambda$
@@
$Sp_{\mathcal{E}/\tilde{f}_{X}}(q^{-1}(\mathbb{V}_{S}^{p}(\lambda)))=\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o}),$
@@
  • $\mathbb{V}_{S}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$, $\mathbb{V}_{S}(\tilde{\lambda})\subset E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$, are the flat leaves, e.g. $\mathbb{V}_{S}(\lambda):=\pi_{S}(\lambda\times\tilde{S}^{an}_{\mathbb{C}})$ where $\pi_{S}:H^{2p}(E^{an},\mathbb{C})\times\tilde{S}_{\mathbb{C}}^{an}\to E^{2p,an}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$ is the morphism induced by the universal covering $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$,
@@
  • Consider now the case where $Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\neq 0$. We have,
    $Sp_{\mathcal{E}/\tilde{f}_{X}}(\mathbb{V}_{S}^{p}(\tilde{\lambda}))=\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))=\overline{\mathbb{V}_{T^{o}}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))}\cap p^{-1}(S_{\mathbb{C}})\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})$
    where $p:=p\otimes\mathbb{C}:E^{2p}_{DR,V_{0}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})\to\bar{T}^{o}_{\mathbb{C}}$ is the projection (the first equality is obvious since $Sp_{\mathcal{E}/\tilde{f}_{X}}$ is a morphism of variation of mixed Hodge structures over $S$, whereas for the second equality the inclusion $\supset$ follows from [6] lemma 2.11 which state the invariance of $\tilde{\lambda}$ under the monodromy
@@
    is defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes. Hence,
    $\mathbb{V}_{S}^{p}(\lambda)=q(q^{-1}(\mathbb{V}_{S}^{p}(\lambda)))=q(\mathbb{V}_{S}^{p}(\tilde{\lambda}))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$
    is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes, where the second equality follows from the fact that $Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\neq 0$ and the fact that $Sp_{\mathcal{E}/\tilde{f}_{X}}(W_{2p-1}E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}}))=0$.
@@
Since $p\in\mathbb{Z}$ and $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$ are arbitrary, the Hodge conjecture holds for $E_{0}$. Proposition 1(ii) and the fact that the Hodge conjecture is true for smooth projective varieties of dimension less or equal to $d-1$ by induction hypothesis then implies that Hodge conjecture holds for $X$, since $X$ is birational to $E_{0}$ and $X,E_{0}\in\mathop{\rm PSmVar}\nolimits(\mathbb{C})$ are connected of dimension $d$. ∎

### v11

@@ 3 Hodge conjecture for smooth projective varieties
$f:\mathcal{Y}^{\prime}_{\mathbb{C}}=V(\tilde{f})\subset\mathbb{P}^{d+1}_{\mathbb{C}}\times\mathbb{P}(l_{X})\to\mathbb{P}(l_{X})$
the universal family of projective hypersurfaces of degree $l_{X}$, and $0_{X}\in\mathbb{P}(l_{X})$ the point corresponding to $Y_{0_{X}}$. Denote $\Delta\subset\mathbb{P}(l_{X})$ the discriminant locus parametrizing the hypersurfaces which are singular. Let $\bar{S}:=\bar{0_{X}}^{\bar{\mathbb{Q}}}\subset\mathbb{P}(l_{X})$ be the $\bar{\mathbb{Q}}$-Zariski closure of $0_{X}$ inside $\mathbb{P}(l_{X})$. If $Y_{0_{X}}$ is smooth, then $X$ satisfy the hodge conjecture since $Y_{0_{X}}$ satisfy the Hodge conjecture ([5]) and by proposition 1(ii) and the induction hypothesis since $X$ is birational to $Y_{0_{X}}$. Hence, we may assume that $Y_{0_{X}}$ is singular, that is $0_{X}\in\Delta$. Then, $\bar{S}\subset\Delta$ since $\Delta\subset\mathbb{P}(l_{X})$ is defined over $\bar{\mathbb{Q}}$. Let $\bar{T}\subset\mathbb{P}(l_{X})$ be an irreducible closed subvariety of dimension $\mathop{\rm dim}\nolimits(\bar{S})+1$ defined over $\bar{\mathbb{Q}}$ such that
@@
We take for $\bar{T}=V(f_{1},\cdots,f_{s})\subset\mathbb{P}(l_{X})$ an irreducible complete intersection defined over $\bar{\mathbb{Q}}$ which intersect $\Delta$ properly and such that
  • $V(f_{1})\supset\mathop{\rm sing}\nolimits(\Delta)\cup\bar{S}$
  • for each $r\in[1,\cdots,s]$, $V(f_{1},\cdots,f_{r})\supset\mathop{\rm sing}\nolimits(V(f_{1},\cdots,f_{r-1})\cap\Delta)\cup\bar{S}$ contain inductively the singular locus of the intersection with $\Delta$ and $\bar{S}$.
@@
is an isomorphism if $\mathop{\rm dim}\nolimits\bar{T}\geq 2$ and surjective if $\mathop{\rm dim}\nolimits\bar{T}=1$. We have then $\bar{T}\cap\Delta=\bar{S}\cup\bar{S}^{\prime}$, where $\bar{S}^{\prime}\subset\bar{T}$ is a divisor. We then consider the family of algebraic varieties defined over $\mathbb{Q}$
@@
  • $t\in S(\mathbb{C})$ is the point such that $\mathcal{Y}^{\prime}_{\mathbb{C},t}=Y_{0_{X}}$ and $\epsilon_{t}:E_{0}:=\mathcal{E}_{0,\mathbb{C},t}\to Y_{0_{X}}$ is surjective (there is at least one component of $E$ dominant over $Y_{O_{X}}$ since $E$ is dominant over $Y_{O_{X}}$ as $\mathcal{E}$ is dominant over $\mathcal{Y}^{\prime}_{\bar{S}}$), in particular $E_{0}$ is birational to $Y_{0_{X}}$,hence $E_{0}$ is birational to $X$. We denote $E_{i}:=\mathcal{E}_{i,\mathbb{C},t}$ for each $0\leq i\leq r$ and $E:=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$.
@@
which is an algebraic variety by [6] (finite over $S$). Note that $\mathbb{V}^{p}_{S}(\lambda)\subset HL^{p,2p}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$ is the union of the irreducible components passing through $\lambda$. Similarly, for $\eta\in H^{2p}(\mathcal{E}_{\mathbb{C},t^{\prime}}^{an},\mathbb{Q})$, $t^{\prime}\in S(\mathbb{C})$, we consider $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$ the universal covering in $\mathop{\rm AnSm}\nolimits(\mathbb{C})$ and
@@
$(E^{2p}_{DR}(\mathcal{Y}_{T^{o}}/T^{o}),F)=H^{2p}\tilde{f}_{X*}E_{zar}(\Omega^{\bullet}_{\mathcal{Y}_{T^{o}}/T^{o}},F_{b})\in\mathop{\rm Vect}\nolimits_{\mathcal{D}fil}(T^{o})$
@@
$(E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o}}/T^{o}),F):=H^{2p}\tilde{f}_{X*}E_{zar}(\Omega^{\bullet}_{\mathcal{Y}_{T^{o}}/T^{o}}(\mathop{\rm log}\nolimits\mathcal{E}),F_{b})\in\mathop{\rm Vect}\nolimits_{\mathcal{D}fil}(\bar{T}^{o})$
@@
$\alpha\in H^{2p}(\mathcal{Y}_{t^{\prime\prime}}^{an},\mathbb{Q})=i_{t^{\prime}}^{*}\psi_{S}R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}_{T^{o},\mathbb{C}}^{an}},\,t^{\prime}\in S_{\mathbb{C}},\,t^{\prime\prime}\in T^{o}(\mathbb{C}),$
@@
  • $\pi_{T^{o}}:\tilde{T}^{o,an}\to T_{\mathbb{C}}^{o,an}$ the universal covering in $\mathop{\rm AnSm}\nolimits(\mathbb{C})$, the factorization
    $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to\tilde{\bar{T}}_{\mathbb{C}}^{o,an}\xrightarrow{\pi_{\bar{T}^{o}}}\bar{T}_{\mathbb{C}}^{o,an}$
    and
@@
    which is an algebraic subvariety (finite over $T^{o}$) by [6],
@@
    $\mathbb{V}_{S}(\alpha):=\pi_{S}(\alpha\times\tilde{S}^{an}_{\mathbb{C}})\subset H^{2p}\psi_{S}E_{DR}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})^{an}=i_{S}^{*mod}(E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o}))^{an}$
@@
where $HL^{2p,p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$, $HL^{2p,p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})$ are the locus of Hodge classes. Let $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$, where we recall $E_{0}=\mathcal{E}_{0,\mathbb{C},t}$ and $E=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$. Consider then $\tilde{\lambda}\in H^{2p}(E^{an},\mathbb{Q})$, such that $q(\tilde{\lambda})=\lambda$, and
$Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\in Sp_{E/f_{X}}(H^{2p}(E^{an},\mathbb{Q}))\subset i_{t}^{*}\psi_{S}R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}_{T^{o},\mathbb{C}}^{an}}.$
@@
using the fact that there exists a neighborhood $V_{\bar{S}\cup\bar{S}^{\prime}}\subset\bar{T}_{\mathbb{C}}$ of $\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}$ in $\bar{T}_{\mathbb{C}}$ for the usual complex topology such that the inclusion $i_{\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}}:\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}\hookrightarrow V_{\bar{S}\cup\bar{S}^{\prime}}$ admits a retraction $r:V_{\bar{S}\cup\bar{S}^{\prime}}\to\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}$ which is an homotopy equivalence. The locus of Hodge classes passing through $\lambda$
@@
$\mathbb{V}_{S}^{p}(\lambda)=q(Sp_{\mathcal{E}/\tilde{f}_{X}}^{-1}(\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}}),$ (2)
@@
  • $\mathbb{V}_{S}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$, $\mathbb{V}_{S}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o})$, are the flat leaves, e.g. $\mathbb{V}_{S}(\lambda):=\pi_{S}(\lambda\times\tilde{S}^{an}_{\mathbb{C}})$ where $\pi_{S}:H^{2p}(E^{an},\mathbb{C})\times\tilde{S}_{\mathbb{C}}^{an}\to E^{2p,an}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$ is the morphism induced by the universal covering $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$,
@@
  • $Sp_{\mathcal{E}/\tilde{f}_{X}}:=Sp_{\mathcal{E}/\tilde{f}_{X}}\otimes\mathbb{C}:E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\to i_{S}^{*mod}E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o})$.
@@
  • Consider now the case where $Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\neq 0$. We have the key equality
    $\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))=\overline{\mathbb{V}_{T^{o}}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))}\cap p^{-1}(S_{\mathbb{C}})\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})$
    where $p:=p\otimes\mathbb{C}:E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})\to\bar{T}^{o}_{\mathbb{C}}$ is the projection, the inclusion $\subset$ is obvious whereas the inclusion $\supset$ follows from [6] lemma 2.11 which state the invariance of $\tilde{\lambda}$ under the monodromy
@@
    is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes. Hence,
    $\mathbb{V}_{S}^{p}(\lambda)=q(Sp_{\mathcal{E}/\tilde{f}_{X}}^{-1}(\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$
    is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes.
@@
Since $p\in\mathbb{Z}$ and $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$ are arbitrary, the Hodge conjecture holds for $E_{0}$. Proposition 1(ii) and the fact that the Hodge conjecture is true for smooth complex projective varieties of dimension less or equal to $d-1$ by induction hypothesis then implies that Hodge conjecture holds for $X$, since $X$ is birational to $E_{0}$ and $X,E_{0}\in\mathop{\rm PSmVar}\nolimits(\mathbb{C})$ are connected of dimension $d$. ∎